BY THE WAY - THIS IS DEADLY SERIOUS AND NO JOKETo qualify for PIP, the DWP must assess that any condition must affect you for more than 50% of the time or more than 3 out of 7 days.
A Judge has ruled that the DWP must consider the probability of separate conditions COMBINING to affect you for more than 50% of the time or 4 or more days a week.
Here is the maths I have just sent to DWP today.
(2) REGARDING LEGALITY – FAILED TO FOLLOW YOUR
OWN GUIDELINES RELATING TO THE PROBABILITY OF MY WIFE SUFFERING FROM THE EFFECTS
OF MULTIPLE CONDITIONS FOR MORE THAN 3 DAYS A WEEK – 50% of the time.
Your decision to award my wife ‘0’ points
after the first assessment and then again after the mandatory reconsideration
is contrary to your regulations and, therefore, incorrect and unlawful.
Your regulations clearly state that a claimants disabilities must affect her
for more than 50% of the time before they can receive PIP.
In order to do this, you must apply the mathematical rules of probability in
relation to the probability of multiple
health conditions/disabilities affecting claimants in any one week (7 days) for
more than 50% of the time.
A judge recently made the following calculations based on a claimant suffering
from two health conditions/disabilities affecting them 3 out of the 7 days of
the week, both, therefore, not qualifying for PIP individually.
_____________________________________________________________________________________
The claimant who had both chronic obstructive pulmonary disease (“COPD”) and
rheumatoid arthritis.
The first tier tribunal found that the claimant had 3 bad days a week
for COPD, and 2 to 3 bad days a week, depending on the weather, for rheumatoid
arthritis. Each condition affected the claimant’s ability to carry out some of
the activities in the PIP test, such as preparing food or dressing.
The tribunal decided that the claimant’s conditions did not cause him to
satisfy any point scoring descriptors for over 50% of the time, as required by
regulation 7, so he was not eligible for PIP.
Probability
theory
The upper tribunal judge, however, held that the
tribunal had failed to find out if the effect of one condition was enough to
cause problems or whether there was only an issue when both conditions applied
at the same time.
If the effects of either condition on its own was enough to allow points
to be scored, it would still be necessary to work out how many days a week on
average, the claimant was affected by one or both conditions. If 50% or less
then no award of PIP could be made.
Relying on a submission by the DWP, the tribunal judge went on to set
out in great detail how to calculate the probability of the two conditions
occurring for more than 50% of the time.
The judge argued:
If we say that the rheumatoid arthritis occurs 3 times
a week then the probability that regulation 7 is satisfied is worked out as
follows:
Probability(A or
= Probability(A) +
Probability(B) – Probability(A and 
The probability of event A (arthritis) occurring,
3/7, is added to the probability of event B (COPD) occurring. 3/7 + 3/7 =6/7.
However, this probability (6/7) also includes the
times on which the COPD and arthritis both occur. This must be subtracted. It
is double counting.
The probability of the COPD and arthritis both
occurring is 3/7 x 3/7 = 9/49.
We need to subtract 9/49 from
6/7. To do this we convert 6/7 into 42/49 (by multiplying by 7). We can
then do the simple sum 42/49
- 9/49 = 33/49. This is the probability of an operative
condition being present on a given day.
33/49 can otherwise be understood as 4.71/7, if you
were looking at this on a weekly basis (as we have here).
Therefore the descriptor is satisfied for more than
50% of the days.
This finding was accepted and is a
precedent to any other cases challenging the effects of multiple health
conditions/disabilities.
__________________________________________________________________________________
Now, expanding this ruling to apply to my wife's several health conditions/disabilities
The probability that at least one condition is suffered on a day is 1
minus the probability that no conditions are suffered. That latter is the
product of the probabilities of not suffering each condition.
Using the example probabilities of 3/7 for suffering the two conditions,
we get 1 - (4/7)*(4/7) = 1 - 16/49 = 33/49.
Extending
this for the conditions my wife suffers from: PTSD, CONFUSION & ANXIETY,
COPD (Shortness of Breath), FYBROMYALGIA, DYSEXECUTIVE SYNDROME (Brain Damage),
SEIZURES, ACUTE CORONARY SYNDROME (Unstable Angina), IBS (Irritable Bowel
Syndrome), BRAIN TUMOUR (Meningioma), LOSS OF BALANCE, ASTHMA.
That totals
10 health conditions/disabilities.
Being generous and saying that there might
some overlapping in the mental health/[psychological problems/disabilities, I will
knock that down to a total of 7 health conditions/disabilities and, as you have
argued (though we vehemently disagree), each one individually affects Angela for
less than 50% of the time, we will say that they affect her 3 days out of 7.
Therefore,
mathematically, each condition affects her 3/7, 3/7, 3/7, 3/7, 3/7, 3/7 and 3/7.
The
probability of any occurrence is 1 - (4/7)*(4/7)*(4/7)*(4/7)*(4/7)*(4/7)*(4/7) = 1 – 16,384/823,543 = 0.98.
Therefore,
the probability of Angela suffering from her multiple health conditions/disabilities in
any one week, is 98% of the time.
Please re-adjust
your PIP assessment to feature in the probability of Angela suffering from her multiple
multiple health
conditions/disabilities for 98% of the time.
NOTE:If you suffer from mulltiple health conditions, you should factor in this argument at any stage in your claim, appeal or new claim.
Smile (it is sorta funny - but very real!)
gjaggers